OpenStudy (anonymous):

(√(2) + √(6)) / √(2 + √(3)) The calculated answer is 2. How would I prove it without using a calculator?

OpenStudy (anonymous):

look how sqrt(3) can also be sqrt(300/100), then break that up to get sqrt(300)/sqrt(100) which will be sqrt(300)/10. So, from there, notice that 17^2 = 289 and 18^2 = 324. So, 17.5 falls between the two. So, sqrt(300) = approximately 17.5. Still keep in mind that the denominator is 10. So you have 17.5 / 10 which is 1.75. Follow this pattern for the rest of them. I hope this makes sense (:

OpenStudy (anonymous):

\[\sqrt{2} + \sqrt{3}\] Is that right?

OpenStudy (anonymous):

\[\sqrt{2+\sqrt{3}}\] Or is it this?

OpenStudy (anonymous):

Sorry for the confusion. It's the second one.

OpenStudy (anonymous):

Top like that as well?

OpenStudy (anonymous):

Let me just rewrite that :) \[(\sqrt{2} + \sqrt{6}) / \sqrt{2 + \sqrt{3}}\]

OpenStudy (anonymous):

k, let me look...

OpenStudy (anonymous):

Well, I haven't done it but there is a standard sort of approach to these things.. U want to get rid of radicals on the bottom where u can so start off by multiplying top and bottom by the bottom to get rid of the first (outer) square root

OpenStudy (anonymous):

Then you have 2 + sqrt 3 only on the bottom and to get rid of that square root multiply by 2 - sqrt 3.

OpenStudy (anonymous):

That brings the bottom to 1.

OpenStudy (anonymous):

Sorry, busy... With me so far...?

OpenStudy (anonymous):

You should just have a numerator now composed of three pieces including the ugly one, right?

OpenStudy (anonymous):

Anybody there?

OpenStudy (anonymous):

I'm still here :) Just trying to make sense of the numerator...

OpenStudy (anonymous):

Now I think we should square it so we can deal with the ugly bit and we will square root it again after...

OpenStudy (anonymous):

Squaring the 1 on the bottom doesn't change it.

OpenStudy (anonymous):

U need to multiply it all out...

OpenStudy (anonymous):

And simplify.

OpenStudy (anonymous):

I'm sorry; you lost me there. so the numerator is made up of \[(\sqrt{2} + \sqrt{6}) (\sqrt{2 + \sqrt{3}}) ( 2 - \sqrt{3)}\] right? and I multiply it out and simplify.

OpenStudy (anonymous):

U can't multiply it like that because of the ugly bit so you have to square the whole thing in order to simplify it and then after it is simplified remember to take the aqrt to get back to where u started. Obviously we want it to come out to 4 and guess what, it does:-)

OpenStudy (anonymous):

What a mess, it was easier with a calculator.....

OpenStudy (anonymous):

U get it yet? Page full of multiplying out and simplifying....

OpenStudy (anonymous):

Still trying to work it out...

OpenStudy (anonymous):

Anyway, you know how to do these in the future, right?

OpenStudy (anonymous):

Yes, thanks for the help :)

OpenStudy (anonymous):

ur welcome;-)

OpenStudy (anonymous):

It is much less work if you just square the thing first rather than later:-( Then it just drops out easily in a couple lines, I should have looked at it abit more closely before jumping in, my apologies. Still, good practice with radicals:-)

OpenStudy (anonymous):

haha ok thanks :)